

********************************************************************
UARS WEEKLY STATUS REPORT - 16 MAR 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is currently in day only operations in support of low spacecraft
power conditions. This two day period started on Thursday and completes at
the end of the day today (Friday). Otherwise, SOLSTICE continues to operate
normally, taking solar data on all available orbital tracking periods and
stellar calibration observations whenever possible. We continue to gather
enough data to meet all the SOLSTICE science objectives during this period
of reduced tape recorder operations.
***********
SUSIM
***********
On Wednesday 14 March 2001, SUSIM went through a planned shutdown to
make more power available during the YAW around. Otherwise, SUSIM
performed its daily solar scans of the sun this past week during periods
of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS was activated in its half power mode on March 19. At the current
time, AXIS is not generating energy deposition. SwRI is looking into this.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
PEM has examined its processed files of the rehost machine and has found
differences between those and the cataloged files. PEM is investigating.
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. PEM has requested DAP command capability with
the CMS from PEM2K::. Tests of the NRT data transfer from CDHF2:: through
PEM2K:: to the unix computer running the PEM GSE software show no difference
between equivalent NRT files transferred from PEM::. Thus, NRT data from
CDHF2:: is captured correctly on PEM2K::. Tests of the NRT data transfer
from CDHFB:: were not successful from both PEM:: and PEM2K::. Results of the
test of NRT data transfers from the rehost computer acted like the rehost
computer did not recognize either PEM machine. Investigation is underway.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 19 April 2001.
***********
CDHF
***********
STATUS
----------
o The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
**45 days SUS (7.2) PROD (68R-67C=1L)
o DAAC DATA:
- Data created in the month of January 2001 has been previewed for
HALOE, NMC, and UKMO in preparation for loading into the DAAC Tables.
o UARS CDHF REHOST:
- The following software has been recompiled and relinked and PC is
testing new executables. Status follows:
- Attitude Extraction V7.3 successful and promoted to production mode.
Being processed on current and rehost systems.
- Sensor Extraction V5.0 successful and promoted to production mode.
Being processed on Rehost System only per FDF.
- Observatory Log V3.4 successful and promoted to production mode.
- HALOE V4.82 successful. HALOE has successfully validated the output
product and has requested an older UARS DAY to process to be compared
with day processed on current system.
- HRDI V4.4 successful. Currently waiting validation results from PI.
- GPI V4.05 successful. Currently waiting validation results from PI.
- PEM V3.32 re-compiled and re-linked.
PEM group still validating the output product created from configured
software dated 6-dec-2000.
- SOLSTICE V4.62 (JOB 1) tested successfully. Currently waiting
validation results from PI.
- SUSIM V7.2 was successful against days designated by the SUSIM PI.
- WINDII V5.11a was installed and configured on the rehost system.
WINDII has been powered off and is not processing against current
data. An older day will be selected and files will be requested
to be migrated.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 March 2001 (Orbit 51796, GMT Day 062) through 09 March
2001 (Orbit 51900, GMT Day 068).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 068/0035 51886 AXIS Power ON (7w)
PEM 068/1156 51893 AXIS 2 OFF
The beta angle ranged between the angles of 43.0 degrees to 25.2 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 2 degrees on 15 March 2001 (DOY 074).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 and 1.0 degrees C. Additional battery performance data for each
battery on 03 March 2001 and 09 March 2001 are provided in Appendix D.
The clock error ranged from Ð11.3 to 10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 23
¥ PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 03
¥ SOLSTICE = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 02/026 - 2287. 49920005
GROUND SYSTEM OPERATION
Two (4) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/067 show the observatory data loss to be 300 hours, 19
minutes, 59.848 seconds. This shows no change since the last report period.
This is a 0.39 percent data loss, which equals a 99.61 percent data capture for
the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS is using both the Fine Sun Sensor (FSS) and the Three Axis Magnetometer
(TAM) as inputs to UFLTR this week. The FOT is pleased with their performance,
but continues to make adjustments to the TAM alignment to improve its
performance and UARSÕ attitude. The FOT awaits the completion of the FSW patch
that will improve the magnetic field model from a 5th order to 7th order
solution.
MEETINGS
The FOT attended a UARS TAM Patch Code Review on Monday, March 05, 2001.
The FOT attended a UARS Monthly on Wednesday, March 07, 2001.
The FOT attended a UARS 3 TDRS Meeting on Friday, March 09, 2001
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 15 March 2001 (DOY 074). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 March 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51799 062/0631 01-135 NBTR NOOPS D-062
51800 062/0815 01-135 NBTR NBTR PBK / RTS-58
51807 062/1832 01-136 NBTR NOOPS D-063
51814 063/0632 01-136 NBTR NOOPS D-063
51815 063/0713 01-136 NBTR NBTR PBK / RTS-58
51822 063/1833 01-137 NBTR NOOPS D-064
51823 063/2018 01-138 OBC Clk Adj 0.00 to 20.6 msec/day
51829 064/0634 01-137 NBTR NOOPS D-064
51830 064/0810 01-137 NBTR NBTR PBK / RTS-58
51837 064/1915 01-139 NBTR NOOPS D-065
51838 064/2021 01-140 MACS TAM ALIGNMENT
51843 065/0355 01-141 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
51844 065/0635 01-139 NBTR NOOPS D-065
51845 065/0822 01-139 NBTR NBTR PBK / RTS-58
51850 065/1555 01-142 MACS TAM ALIGNMENT
51852 065/1836 01-143 NBTR NOOPS D-066
51859 066/0637 01-143 NBTR NOOPS D-066
51861 066/0825 01-143 NBTR NBTR PBK / NBTR-B TO REC.
51867 066/1838 01-144 NBTR NOOPS D-067
51880 067/1557 01-145 MACS TAM ALIGNMENT
51882 067/1918 01-146 NBTR NOOPS D-068
51886 068/0035 01-147 PEM AXIS Power ON (7w)
51893 068/1156 01-148 PEM AXIS 2 OFF
51896 068/1738 01-146 NBTR NOOPS D068
51897 068/1843 01-149 NBTR NOOPS D069
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 March 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
SOMECRLDNGXA SOLS CMD ERROR 51883
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 March 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51823 063/2018 Clk Adj 0.00 to 20.6 msec/day
51843 065/0355 Clk Adj 20.6 to 0.00 msec/day
_____________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 March 2001
03 Mar 2001 - GMT Day 062 Beta = 43.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- --------- ----- ------- -----
1 +016.8/+00.0 -3.68/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-44.8 +5.15/+3.87 +26.8/-15.2 25.1 1.030 20.5
3 +22.4/-67.2 +4.51/+2.92 +26.4/-15.2 25.1 1.020 20.2
09 Mar 2001 - GMT Day 068 Beta = 25.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- --------- ------ ------- -----
1 +016.8/+00.0 -3.05/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +16.8/-89.6 +5.79/+4.51 +30.0/-14.4 24.5 1.030 21.1
3 +22.4/-95.2 +4.51/+3.55 +29.6/-14.4 24.3 1.020 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_____________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 March 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT 51803 061/0206 14170 Generic Late Acquisition
01-014 FOT 51820 061/0206 14170 Generic Late Acquisition
01-015 FOT 51839 064/2245 23697 Late Acq. and Dropout
01-016 SOLSTICE 51883 067/2029 N/A SOLS CMD ERR/NO DATA LOSS
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-015 FOT 51839 064/2245 23697 Late Acq. and Dropout
01-016 SOLSTICE 51883 067/2029 N/A SOLS CMD ERR/NO DATA LOSS
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT see list Mar 01 14170 Generic Late Acquisition
SEVNTFW01/10